OpenStudy (arbyj):

can someone correct me if i set up this problem incorrectly? (it'll be in the comments)

OpenStudy (arbyj):



OpenStudy (anonymous):

yes you have set it up correctly. Well that's what I believe.

OpenStudy (aum):

10^6 micrograms = 1 gram 48.2 micrograms = 48.2 / 10^6 = 48.2 x 10^-6 = 4.82 x 10 x 10^-6 = 4.82 x 10^-5 gram. 10^3 grams = 1 Kg 4.82 x 10^-5 grams = 4.82 x 10^-5 / 10^3 = 4.82 x 10^-5 x 10^-3 = 4.82 x 10^-8 Kg.